OpenStudy (anonymous):

Solve 6 sqrt(x+1) A. x = 8 B. x = 71 C. x = 75 D. x = 107

OpenStudy (anonymous):

\[6\sqrt{x+1}\]

OpenStudy (anonymous):

what's the equation? you can't solve x with an equation.

OpenStudy (anonymous):

without*

OpenStudy (anonymous):

based off of the answer choices I am guessing we are solving for x

OpenStudy (anonymous):

an equation has an = sign it would look something like: 6sqrt(x+1)=0 6sqrt(x+1) itself cannot be solved for x

OpenStudy (freckles):

@AydenCarpenter14 can you copy the problem exactly as you see it?

OpenStudy (freckles):

Use the draw tool.

OpenStudy (freckles):

Again this is all we see in your post "Solve 6 sqrt(x+1) A. x = 8 B. x = 71 C. x = 75 D. x = 107 " There should be an equation. Perhaps 6 sqrt(x+1)=18 or something else

OpenStudy (anonymous):

@freckles @Auxuris Here is a snapshot of the question

OpenStudy (freckles):

This question is incomplete then.

OpenStudy (anonymous):

impossible error question but if you want me to make an educated guess I'd guess A it's the only option that allows a perfect square

OpenStudy (freckles):

Teachers make mistakes sometimes when making tests. It happens.

OpenStudy (anonymous):

Solve 6the square root of x plus one) - 3 = 15</div> <div class= that's what happens when I copy and paste the question?

OpenStudy (freckles):

\[6\sqrt{x+1}-3=15 ?\]

OpenStudy (freckles):

this is the question?

OpenStudy (freckles):

This is an equation and can be solved.

OpenStudy (anonymous):